Venture capittal. Henry invested $12,00 in a new resturant.When the resturant was sold two years laters, he received $27,000. Find his average annual return by solving the equation 12,000(1+r)^2= 27,ooo.
Answer by rfer(16322)

(1+r)^2=27000/12000
(1+r)^2=2.25
1+r=1.5
r=.5 or 50%